OpenStudy (anonymous):

$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155. How much money was invested at 3% interest?

OpenStudy (anonymous):

is this consumer math?

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

0.03x+0.035y = 155 x + y = 5000 x = 5000-y Plug it into the first one 0.03(5000-y) + 0.035y = 155 150 - 0.03y + 0.035y = 155 0.005y = 5 y = 1000 <~~Answer that was put into the 3.5% account. Since 1000 was put into the 3.5% account, 4000 was put into the 3% account. Answer: $4000